# A Retrospective Systematic Study on Hierarchical Sparse Query Transformer-assisted Ultrasound Screening for Early Hepatocellular Carcinoma

### v1

@@ 4.2.1 SOTA Model Comparison
Transitioning to the Multi-Center Test, which evaluates model generalizability across different institutions, HSQformer-B achieves leading scores in Accuracy (88.29±0.71%), AUC (95.38±0.33%), and Recall (90.38±4.13%). This suggests that HSQformer not only excels in identifying positive cases but also maintains high consistency across diverse datasets. In comparison, ConvNext exhibits strong overall performance, particularly in Accuracy (86.39±1.11%) and AUC (94.45±0.53%), yet falls short when contrasted against HSQformer’s results.
@@
For the High-Risk Patient Test, aimed at assessing the model’s effectiveness on patients deemed high-risk, HSQformer-B again leads with top-tier metrics: Accuracy (85.41±0.58%), AUC (88.32±0.59%), and F1 score (90.69±0.38%). Notably, the Recall of 94.09±2.75% underscores its proficiency in detecting critical cases. ViT, while performing well with an Accuracy of 84.02±1.35% and Recall of 92.4±2.19%, does not match the comprehensive superiority of HSQformer-B.
@@
Table 3: This table compares the diagnostic performance of radiologists and AI, showing that AI has surpassed the accuracy of junior radiologists and is comparable to the diagnostic level of senior radiologists.
@@
  • HSQformer-B demonstrated significantly higher Recall (94.09%) compared to senior radiologists, with only a slight decrease in Precision (87.62% vs. 91.76%), aligning closely with the Precision of junior radiologists (87.62% vs. 88.62%). This indicates that while HSQformer-B is highly sensitive in detecting malignant cases, it maintains a balance between Recall and Precision comparable to experienced clinicians.
  • HSQformer-B consistently outperformed the average performance of junior radiologists across all metrics and matched the performance of senior radiologists. Notably, HSQformer-B achieved significant improvements over the overall average of all radiologists in terms of F1 score (90.69% vs. 85.59%) and AUC (88.32% vs. 74.45%). These results underscore HSQformer’s practical effectiveness in HCC clinical screening scenarios.
@@
These observations highlight the potential of HSQformer as a powerful tool in aiding clinical decision-making, particularly in high-stakes diagnostic contexts. The model’s ability to surpass the performance of less experienced clinicians and match that of seasoned experts suggests its utility in enhancing diagnostic accuracy and consistency. Furthermore, HSQformer’s robust performance in critical metrics such as Recall and AUC underscores its value in reducing missed diagnoses and improving patient outcomes.
@@ 4.3.1 Macro Ablations
Expanding on the findings depicted in Figure 4, this research delves into the influence exerted by the proportional parameters at various stages on the model’s performance. Employing state-of-the-art (SOTA) models as a reference point within the field[16, 53, 25, 56, 57], we conducted a comprehensive assessment of the model’s performance under multiple proportional settings, as elucidated in Figure 5. The results indicate that the HSQformer model demonstrates a negligible variation in performance with respect to different stage ratios, particularly with regard to the AUC, which shows a minimal trend of change. Despite this robustness, the experimental data has guided us in selecting the most effective ratio configuration, which is 2:2:6:2, to optimize the model’s performance.
@@
Within the framework of model architecture, the parameters of Query number and dimension are crucial determinants of the model’s efficiency and efficacy. The Query number delineates the number of tokens the model queries during the information processing, signifying the model’s capability to manage a specific number of tokens in a single pass. The Query dimension, conversely, pertains to the dimension associated with these queries, which fundamentally establishes the capacity of each query to capture information. In the Figure 6, we meticulously examined the impact of query number and dimension on model performance. Specifically, when query quantity was held constant, incrementally increasing query dimension significantly enhanced the model’s F1 score and AUC. For instance, with a query number of 50, the F1 score improved from approximately 86.12% to about 87.17%, and the AUC score increased from approximately 94.04% to around 94.74% as the query dimension was augmented from 96 to 768. This suggests that a higher query dimension facilitates the model’s ability to capture long-range dependencies and subtle features, thereby enhancing classification performance.
Further analysis revealed that, with the query dimension held constant, a moderate increase in the number of queries significantly enhanced model performance. Specifically, at a query dimension of 768, the F1 score improved from approximately 87.13% to approximately 88.53%, and the AUC score increased from approximately 94.74% to approximately 95.09% as the number of queries increased from 50 to 100. However, when the number of queries was further increased to 300 and 400, the model performance, while remaining high, did not exhibit significant growth, a similar observation made at a dimension of 96. Moreover, concurrently increasing the number of queries and dimension can markedly improve model performance; for example, raising the number of queries and dimension from 50 and 96 to 400 and 768, respectively, resulted in an F1 score increase from approximately 86.12% to approximately 88.99%, and an AUC increase from approximately 94.04% to approximately 95.22%. Beyond a certain threshold, however, the performance enhancement plateaued, indicating the existence of an optimal balance point. It is important to highlight the trade-off between the number of queries and their dimension. Incrementing either parameter escalates computational costs and amplifies model complexity. Consequently, identifying the optimal equilibrium between these two parameters is pivotal for attaining superior performance while preserving computational efficiency.
@@ 4.3.2 Micro Ablations
This section primarily investigates the efficacy of the CSM at a micro-level design. We commence by examining the sequence of self-attention and cross-attention mechanisms, as depicted in Figure 7. Through our analysis, we ascertained that there is a negligible difference between the serial and parallel approaches. Nonetheless, we opted for the serial approach based on the hypothesis that query tokens initially capture relevant features through cross-attention, subsequently reinforcing these features via self-attention. This sequential methodology aligns with the design philosophy of the Transformer architecture, ensuring the rationality of our design.
@@
Figure 7: Serial vs. Parallel attention mechanisms: structural diagram and performance comparison.
@@
In the study of Mixture of Experts (MoE) models, the number of experts and the Top-K selection strategy employed by the router play a crucial role in determining model performance. In sparse MoE models, where the aim is to maintain computational efficiency, existing research predominantly adopts Top-1 or Top-2 selection strategies. This approach is taken because selecting more experts could undermine the sparsity advantage and increase computational overhead. Conversely, when all experts are activated irrespective of input, this dense configuration does not prioritize sparsity but instead focuses on leveraging the collective knowledge of all experts for potentially better performance. Figure 8 provides an insightful examination into the specific impacts of varying expert counts and routing strategies on model performance.When the number of experts is limited, MoE does not exhibit performance improvements over traditional MLP, which can be considered as having zero experts. Specifically, with only two experts and a Top-1 routing strategy, MoE achieves AUC (Area Under the Curve) scores that are comparable to those of MLP but does not show any significant improvement in F1 score; in fact, there is a decrease of 0.88% in the F1 score compared to the MLP. Furthermore, no substantial differences are observed between the Top-1 and Top-2 routing strategies under these conditions, indicating that a limited number of experts does not contribute to enhanced model performance. These results collectively suggest that when the number of experts is small($\leq 2$), MoE fails to provide performance benefits over conventional architectures like MLP, highlighting the necessity of increasing the number of experts or exploring alternative strategies to achieve superior performance in MoE. Building on this, as the number of experts in MoE increases to four or eight, significant performance improvements over MLPs are observed with a Top-1 routing strategy. The adoption of a Top-2 routing strategy further amplifies these performance benefits compared to the Top-1 approach. Interestingly, activating all experts in a dense MoE configuration does not lead to additional performance gains; instead, it may cause a slight decline, potentially due to the increased model complexity that could result in overfitting. Notably, under the Top-2 routing strategy, MoE with four and eight experts exhibit similar performance outcomes, indicating that four experts might provide an optimal balance between performance and efficiency. This insight is crucial for the design of efficient and high-performing MoE architectures.
@@ 5 Discussion and Limitations
Our research primarily focuses on evaluating the efficacy of AI technology in HCC screening, comparing its diagnostic performance against that of clinical radiologists. The results indicate that AI surpasses the diagnostic performance of junior radiologists and matches or even exceeds that of senior radiologists under certain conditions. This aligns with recent advancements in medical imaging analysis where AI has demonstrated significant improvements in both the speed and accuracy of diagnoses. By leveraging deep learning algorithms, AI models can analyze vast datasets of medical images to detect subtle patterns and features that may elude human observers, thereby aiding clinical decision-making. The designed AI model exhibited robust generalization capabilities across three distinct scenarios, underscoring its adaptability to variations in HCC imaging due to individual differences, tumor size, location, and growth rate. This consistency is particularly crucial for early-stage HCC screening as timely diagnosis plays a pivotal role in improving patient survival rates.However, while discussing the potential applications of AI in HCC diagnostics, it is imperative to acknowledge the challenges it faces in clinical practice. First, training and validating AI models require substantial amounts of high-quality data, which are time-consuming and costly to collect and annotate. Second, despite our provision of some visual insights into the model predictions, the interpretability of AI models remains an ongoing challenge, particularly when it comes to explaining the decision-making process to patients and healthcare providers.
While this study yields promising outcomes, several limitations must be addressed. The HSQformer, specifically tailored for ultrasonic HCC screening, incorporates hyperparameters such as query number and dimension, and the number of experts and Top-K selection strategy within the Mixture of Experts (MoE) framework. The applicability of these parameters to other domains requires further investigation. Additionally, our study emphasizes the efficacy of the model without delving into its parameter count. Although MoE enhances performance, it also substantially increases model size; however, only a subset of experts is activated during inference. Future work will explore the feasibility of using more compact models without compromising performance to enhance accessibility and broader application.
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grants 12326609, 82030047, 82272076, 82371983 and 82171960.
@@ A Architecture Details
Table 4: Configuration of Small, Base, and Large models.

### v2

@@ 4.2.1 SOTA Model Comparison
Transitioning to the multi-center test, which evaluates model generalizability across different institutions, HSQformer-B achieves leading scores in Accuracy (88.29±0.71%), AUC (95.38±0.33%), and Recall (90.38±4.13%). This suggests that HSQformer not only excels in identifying positive cases but also maintains high consistency across diverse datasets. ConvNext has strong overall performance, particularly in Accuracy (86.39 ± 1.11%) and AUC (94.45 ± 0.53%), but falls short when compared to HSQformer.
@@
In the high-risk patient test, which is designed to evaluate the model’s effectiveness on patients identified as high-risk, HSQformer-B once again achieves top-tier metrics: Accuracy (85.41±0.58%), AUC (88.32±0.59%), and F1 score (90.69±0.38%). Notably, the Recall of 94.09±2.75% underscores its proficiency in detecting critical cases. ViT, while performing well with an Accuracy of 84.02±1.35% and Recall of 92.4±2.19%, does not match the comprehensive superiority of HSQformer-B.
@@
Table 3: Comparison of Radiologists and AI. This table compares the diagnostic performance of radiologists and AI, showing that AI has surpassed the accuracy of junior radiologists and is comparable to the diagnostic level of senior radiologists.
@@
  • HSQformer-B demonstrated a significantly higher Recall (94.09%) than senior radiologists, with only a slight decrease in Precision (87.62% vs. 91.76%). Its Precision aligns closely with juniors’ (87.62% vs. 88.62%), indicating a balance between Recall and Precision.
  • HSQformer-B consistently outperformed junior radiologists across all metrics and matched senior radiologists. Notably, it achieved significant improvements over the overall average of all radiologists in F1 score (90.69% vs. 85.59%) and AUC (88.32% vs. 74.45%).
@@
These findings underscore HSQformer’s potential as a clinical decision-support tool, especially in critical diagnostic scenarios. Its performance exceeds that of juniors and matches seniors, indicating its role in improving diagnostic accuracy and consistency. The model’s strong Recall and AUC metrics highlight its effectiveness in minimizing missed diagnoses and enhancing patient outcomes.
@@ 4.3.1 Macro Ablations
To further investigate the influence of stage ratios on model performance, we used SOTA models[6, 47, 8, 49, 51] as references and conducted extensive assessments under various proportional settings, as shown in Figure 5. The results indicate that the HSQformer model exhibits minimal performance variation across different stage ratios, with the AUC showing particularly slight changes. This suggests a certain level of robustness in the model’s performance regarding stage ratios. However, based on the experimental data, we identified the most effective ratio configuration as 2:2:6:2, which optimizes the model’s performance while balancing computational efficiency and accuracy.
@@
Within the model architecture, Query number and dimension are key parameters determining the model’s efficiency and efficacy. The Query number indicates how many tokens the model can process in one pass, while the Query dimension determines each query’s capacity to capture information.
Figure 6 shows that increasing the Query dimension at a constant Query number significantly boosts the model’s F1 and AUC scores. For example, with 50 queries, the F1 score rises from about 86.12% to 87.17%, and the AUC increases from approximately 94.04% to 94.74% when the Query dimension grows from 96 to 768. This means a higher Query dimension helps the model capture long - range dependencies and subtle features better, improving classification. Conversely, moderately increasing the Query number at a fixed dimension also enhances performance. At a Query dimension of 768, the F1 score goes up from about 87.13% to 88.53%, and the AUC increases from around 94.74% to 95.09% as the Query number rises from 50 to 100. However, further increasing the Query number to 300 or 400 yields little improvement. When both parameters increase concurrently, model performance can rise significantly. For instance, raising the Query number and dimension from 50 and 96 to 400 and 768 respectively, the F1 score jumps from about 86.12% to 88.99%, and the AUC increases from approximately 94.04% to 95.22%. Yet, beyond a certain point, the enhancement plateaus, indicating an optimal balance. It’s crucial to note the trade - off between these parameters. Increasing either raises computational costs and model complexity. Thus, finding the optimal equilibrium is vital for achieving superior performance while maintaining computational efficiency.
@@ 4.3.2 Micro Ablations
This section delves into the micro - level design efficacy of the CSM mechanism. As shown in Figure 7, we analyze the sequence of self - attention and cross - attention. Results reveal minimal performance difference between serial and parallel approaches. However, we choose the serial method. This is because, in this sequence, query tokens first capture relevant features through cross - attention, then self - attention reinforces these features. This sequential approach aligns with the Transformer architecture’s design philosophy, ensuring our design’s rationality.
@@
Figure 7: Serial vs. Parallel Attention Mechanisms: Structural Diagram and Performance Comparison
@@
In the study of Mixture of Experts (MoE) models, the number of experts and the Top-K selection strategy are crucial for model performance. In sparse MoE models aiming for computational efficiency, existing research mostly uses Top-1 or Top-2 strategies. This is because selecting more experts may reduce sparsity and increase computational overhead. In contrast, dense configurations, where all experts are activated regardless of input, focus on leveraging all experts’ collective knowledge for potentially better performance.
Figure 8 examines the impacts of varying expert counts and routing strategies on model performance. When the number of experts is limited, MoE shows no performance improvement over traditional MLP (which can be seen as having zero experts). Specifically, with two experts and a Top-1 strategy, MoE achieves comparable AUC scores to MLP but shows no significant F1 score improvement, with a 0.88% decrease compared to MLP. There are also no substantial differences between Top-1 and Top-2 strategies in this case, indicating that a limited number of experts does not enhance model performance. These results suggest that when the number of experts is small ($\leq 2$), MoE offers no performance benefits over conventional architectures like MLP. This highlights the need to increase the number of experts or explore alternative strategies for superior MoE performance. As the number of MoE experts increases to four or eight, significant performance improvements over MLPs are observed with a Top-1 strategy. Using a Top-2 strategy further enhances these benefits compared to Top-1. Interestingly, activating all experts in a dense MoE configuration does not lead to additional performance gains and may even cause a slight decline, potentially due to increased model complexity and overfitting. Notably, under the Top-2 strategy, MoE with four and eight experts shows similar performance, indicating that four experts might provide an optimal balance between performance and efficiency. This insight is crucial for designing efficient and high-performing MoE architectures.
@@ 5 Discussion and Limitations
Our research centers on evaluating AI’s efficacy in HCC screening and comparing its diagnostic performance with that of clinical radiologists. The results reveal that AI outperforms junior radiologists and matches or even surpasses senior ones under specific conditions. This is consistent with recent progress in medical imaging analysis, where AI has significantly enhanced diagnostic speed and accuracy. The designed AI model shows robust generalization across three distinct scenarios, highlighting its adaptability to variations in HCC imaging caused by individual differences, tumor size, location, and growth rate. This consistency is especially vital for early-stage HCC screening, where timely diagnosis is crucial for improving patient survival rates. However, when considering AI’s potential applications in HCC diagnostics, it is essential to recognize the challenges in clinical practice. Firstly, training and validating AI models require large amounts of high-quality data, which are time-consuming and costly to collect and annotate. Secondly, despite some visual insights into model predictions, as provided in the appendix, AI models’ interpretability remains a challenge, particularly in explaining the decision-making process to patients and radiologists.
While our study shows promising results, several limitations must be addressed. The HSQformer, tailored for ultrasonic HCC screening, includes hyperparameters such as query number and dimension, and the number of experts and Top-K selection strategy within the MoE framework. The applicability of these parameters to other domains needs further investigation. Additionally, our study focuses on the model’s efficacy without considering its parameter count. Although MoE improves performance, it also increases model size; however, only a subset of experts is activated during inference. Future work will explore using more compact models without compromising performance to enhance accessibility and broader application.
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grants 12326609, 82272076, 82371983 and 82171960.
@@ A Architecture Details
Table 4: Configuration of Small, Base, and Large Models